OpenStudy (anonymous):

3(2x+3(x-4x))-5x+1 Is the answer -44x+1? if not where did i go wrong?

OpenStudy (anonymous):

No, and it's hard to tell where you did wrong without knowing how you did it. Assuming 3(2x+3(x-4x))-5x+1=0

OpenStudy (anonymous):

Well you are suppose to simplify. I Started with x-4x and got -5x then i did 3 multiplied by -5x and got -15x. then i did 2x+-15x and got -13x. then i did 3 multiplied by -13x and got -39x. then i did -39x-5x and got -44x+1

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

3(x-4x)=3x-12x (2x+(3x-12x)= 5x-12x 3*-7x=-21x -21x-5x=-26x -26x+1=0 / 1-26x=0

OpenStudy (anonymous):

How did you get to that first step than from what i had before?

OpenStudy (anonymous):

You got 3(2x+3(x-4x))-5x+1=0 right? Then i solved the inner ( ) first: 3(x-4x)=3x-12x Then the one outside that : (2x+(3x-12x)= 5x-12x Then the last one: 3*-7x=-21x Then what you get that 3(2x+3(x-4x)) equals -21x Then the rest -21x-5x=-26x And ad the 1 at the end: 1-26x=0